[Brahms-l] Shift report 20050223 00:00-08:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Wed Feb 23 2005 - 07:21:06 EST
Shifter leader was CN
Shifters were HHD


1. machine has been good. Have had beam for all the shift except for 30 min. between stores. Shift ended at 0720 due to beam experiments.

2. runs

MRS@40A1050, FS@2.3A3450

runno(tr2,tr3):13402(226k,137k),13403(368k, 194k), 13404(230k, 246k), 13414(360k, 187k)

non-TPC pedestal run: 13415
TPC pedestal run: 13416

3. Minor problem with the DAQ. Solved by restarting the SBE boards and run/io-servers several times.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Wed Feb 23 07:21:26 2005

This archive was generated by hypermail 2.1.8 : Wed Feb 23 2005 - 07:21:29 EST